A rock is thrown horizontally from the top of a cliff 88 m high with a horizontal speed of 25 m/s, what is the direction of velocity?


1
Expert's answer
2022-05-14T16:45:24-0400

"v_x=25~\frac ms,"

"v_y=\sqrt{2gh}=48~\frac ms,"

"v=\sqrt{v_x^2+v_y^2}=53.7~\frac ms,"

"\alpha=-\arctan\frac{v_y}{v_x}=-63.4°."